A metal sphere A of radius ‘a’ is charged to potential V. What will be its potential if it is enclosed by a spherical conducting shell B of radius ‘b” and the two are connected by a wire. ??

Grade:12

3 Answers

Abhishek Kumar
8 years ago
When it is connected by wire, all charge from inside is conducted to the outer shell.
And then potential inside the shell is a constant and equal to the potential at the surface i.e kQ/b
kQ/a = V So, Q = Va/K
So, New Potential = k(Va/k)/b = Va/b
varun
5 years ago
Charges each equal to Q coloumb are placed at the corners of hexagon of side 'a',the potential at the centre of the hexagon
 
Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your question below.
 
When it is connected by wire, all charge from inside is conducted to the outer shell.
And then potential inside the shell is a constant and equal to the potential at the surface i.e kQ/b
kQ/a = V So, Q = Va/K
So, New Potential = k(Va/k)/b = Va/b
 
Hope it helps.
Thanks and regards,
Kushagra